# Satellite Data Shows Resilience of Tigrayan Farmers in Crop Cultivation During Civil War

### v1

@@ 2.3.5 Post-processing
We obtained a small ground-truth survey dataset collected in Tigray during 2021 by University of Ghent and Mekelle University [27]. This dataset contains ground-truth labels of agricultural fields in Tigray with annotations of “cultivated” (127 fields) or “fallow” (34 fields). To better understand the difficulty of distinguishing cultivated from fallow fields (which may resemble cultivated fields due to the presence of weeds or other vegetation) in satellite observations, we plotted the 12-month time series profiles in each of the 12 Sentinel-2 channels used for our model input. Figure 3 shows the NDVI time series and 7 shows the time series for all Sentinel-2 bands. Figure 7 shows that the time series for all bands is visually very similar for the cultivated and fallow fields surveyed. This observation, in addition to visual inspection of predicted crop maps in Tigray, suggested that our predicted crop maps may overestimate cultivated area in Tigray by classifying some fallow fields as cultivated.
@@
Figure 3: Left: 12-month time series of Sentinel-2 NDVI for ground-truth plots annotated as cultivated (“Crop”), fallow with weeds (“Fallow Weed”), or fallow with no vegetation at all (“Fallow None”). Solid lines depict the mean and shaded regions depict the standard deviation of the time series for all samples. Right: Visualizations of example ground-truth plots in PlanetScope September 2021 basemap. The plot labeled “Fallow with weeds or grass” has a similar appearance to fields labeled with crop types, while the field labeled “Fallow (no vegetaion at all)” is easier to distinguish. Note that our map classification used Sentinel-2 inputs (left-most time series) while our response design used PlanetScope data for photointerpretation of reference sample labels (similar to those shown on the right).
@@
In Figure 3, some differentiation between cultivated and fallow fields can be seen in the NDVI time series: the peak of the NDVI time series appears substantially higher on average than the peak reached in fallow fields. This also makes intuitive sense: we would expect fallow fields to have lower biomass than cultivated fields, since the pattern of vegetation in fallow fields should be sparse and unmaintained. Thus, we hypothesized that for a given predicted cropland map, within the pixels predicted as cropland there will be a minority sub-population of pixels that reached lower peak NDVI than the majority of the pixels which may correspond to fallow fields (or other erroneous predictions). To identify and remove these potential false positives, we computed the mean and standard deviation of the peak NDVI value for all positive (cropland) predictions in a given map, and identified which of those pixels had NDVI above or below $n$ standard deviations from the mean (i.e., $\mu\pm n\sigma$). We then re-classified those “anomalous” positive pixels to the non-crop class. We evaluated the effect of this post-processing algorithm using $n\in[0,0.5,1.0,1.5,2.0,2.5,3.0,3.5]$ on the true positive rate (TPR) versus false positive rate (FPR) for the ground-truth dataset and validation dataset for Tigray in 2021. We found that a threshold of $n=3.5$ offered the optimal tradeoff between TPR and FPR for the ground-truth dataset, without significantly increasing the FPR for the validation dataset. Thus for our final cropland maps in 2020 and 2021, we applied this post-processing algorithm with $n=3.5$ in an effort to reduce false positive predictions.
@@ 2.4.1 Sampling and response design
We created a stratified random sample of $10\times 10$-meter pixels using the four strata in the 2020-2021 cropland change map classes. We drew 799 samples total (sampling rate 0.00015% in Tigray) using the following allocation to the four strata: 131 for “stable planted,” 404 for “stable not-planted,” 134 for “planted gain,” and 130 for “planted loss.” We pre-allocated 100 samples to each change stratum and then allocated the remaining samples approximately proportionally to the map area for each strata to reduce the uncertainties in area estimates for these classes resulting from low sample sizes [20]. We labeled reference samples using visual interpretation of satellite imagery by trained individuals, using the same procedure and satellite data sources described in Section 2.3.4. Annotators answered two questions, “Does this point contain active cropland in 2020?” and “Does this point contain active cropland in 2021?”, for each point based on whether the point location visually resembled an agricultural field and showed signs of sowing, growing, and harvesting during the approximate crop calendar of Tigray. Each reference sample was independently labeled by at least two annotators. For samples that did not have unanimous label agreement for both questions, we convened expert annotators as a group to discuss and decide the label to assign for each year.
2.4.2 Cultivated area in Tigray and admin2 zones
@@
We used the reference samples to compute confusion matrices between the map class and reference labels, which were then used to compute map accuracy, area estimates and uncertainties, and investigate changes between 2020-2021 that could be seen visually in satellite images [20]. We computed area estimates for the four change classes representing transitions from 2020 to 2021 as well as annual cropland area estimates for 2020 and 2021 separately. For the annual area estimates, we used the reference labels from each year provided by the two-year change reference samples. We computed the annual and change area estimates for the Tigray region as well as within each of the five admin2 zones in Tigray. For each zone, we clipped the classification map to the zone boundary and filtered the reference samples to use only the samples within a given zone to estimate area. The code used to compute annual area estimates is available at https://github.com/nasaharvest/crop-mask/blob/165d01cc3114427fac59f28ec2eeee5c97de25fc/notebooks/crop_area_estimation.ipynb. The code used to compute change area estimates is available at https://github.com/nasaharvest/crop-mask/blob/359708379fb7f064e1248ff35daaf56e27eee97a/notebooks/ethiopia_tigray_change_area_estimation.ipynb. The reference samples can be accessed via Github following the paths recorded in the change area estimation notebook.
@@ 3.1 Map classification and accuracy assessment
Figure 4: Annual cropland and change maps overlaid with admin2 zone boundaries (solid lines) and 5 km buffer around conflict events (dashed lines).
Table 1 and Table 2 report the overall accuracy, precision (user’s accuracy), recall (producer’s accuracy), and F1 score for the 2020 and 2021 maps, respectively. We computed these metrics using the independent test sets labeled for each year described in Section 2.3.4. We found F1 scores of $0.78\pm 0.11$ for the 2021 map and $0.69\pm 0.11$ for 2020. Overall accuracy was similar between the two years: $0.83\pm 0.02$ in 2020 and $0.84\pm 0.02$ in 2021. There was a moderate difference in precision (user’s accuracy) between 2020 ($0.75\pm 0.04$) and 2021 ($0.79\pm 0.04$) and larger differences in recall (producer’s accuracy) between 2020 ($0.65\pm 0.03$) and 2021 ($0.76\pm 0.03$). This is likely due to differences in climatic factors between the two years that result in different types of errors made by the classifier. Unbalanced precision scores between the two years may result in higher error rates when the maps are combined to form a change map, compared to maps with balanced precision scores.
@@
Table 1: Performance metrics for 2020 test set used for assessing accuracy of annual cropland map. We report metrics for other cropland and land cover maps on the same test set for comparison to our map (ordered by decreasing F1 score).
@@
Table 2: Performance metrics for 2021 test set used for assessing accuracy of annual cropland map. We report metrics for other cropland and land cover maps on the same test set for comparison to our map (ordered by decreasing F1 score).
Some prior studies used existing global land cover maps to derive a cropland map for change assessments instead of creating a custom classification map as in our study; for example, Weldegebriel et al. [19] used the WorldCover 10 m/pixel global land cover map for 2020 [29] to assess changes in cultivated area between 2020-2021 in Tigray. To assess the accuracy of our custom map classification compared to existing land cover maps, we evaluated multiple global land cover maps using our independent test set following the methodology described in Kerner et al. [30]. We only included maps that were created in 2017 or later to ensure temporal coherence: WorldCover [31, 32], Digital Earth Africa [33, 34], Copernicus [35, 36], Dynamic World [37], ASAP [38, 39], GLAD [40, 41], and Esri LULC [42, 43]. If multiple years were provided by a land cover product, we used the closest year to our reference data year. Of the compared maps, we found that WorldCover had the highest performance metrics overall; the F1 score for the WorldCover map in 2020 was comparable to our map but was lower for the 2021 map.
@@
Table 3: Annual area estimates for Tigray and each admin2 zone within Tigray (units are kha)
@@ 3.2 Annual area estimates
We report the annual cropland area estimates for 2020 and 2021 in Table 3. Figure 5 visualizes these estimates as a bar plot to illustrate the relative differences between each year. We found similar estimates for cultivated area in Tigray in 2020 and 2021, with a possible small net increase: $1,132\pm 133$ kha in 2020 and $1,217\pm 132$ kha in 2021 (1 kha = 1,000 ha). We found a similar trend for each admin2 zone, with the exception of the Central zone which showed a possible small net decrease. This net decrease in Central Tigray may be due to a higher concentration of conflict events recorded in the Central zone during the growing season [19]. Figure 6 shows the confusion matrices between the map and reference samples for Tigray. We provide the confusion matrices for admin2 zones in Figures 8-12. All confusion matrices are expressed in terms of proportion of mapped area based on good practice recommendations [20].
Figure 5: Annual area estimates for Tigray and each admin2 zone within Tigray (units are kha)
@@ 3.3 Change area estimates
Table 4 summarizes the area estimates for the four transition classes. We found that cropland gain accounted for between $70-160$ kha while cropland loss accounted for $19-79$ kha in Tigray. The stable transition classes accounted for the majority of the area in Tigray: $1,033-1,293$ kha of stable cropland and $3,798-4,054$ kha of stable non-cropland. Figure 5(a) shows the confusion matrix between the reference samples (stratified by the four change class strata) and the classified change map, expressed in terms of mapped area proportions. For Tigray, there was good agreement between the map and reference samples for the stable not planted class: of the area mapped as stable not planted, only a small fraction of the mapped area should have been mapped as stable planted according to the reference sample. There was also good agreement between the map and reference samples for the stable planted class, though approximately one third of the area mapped as stable planted should have been mapped as stable not planted according to the reference sample. This is likely due to the similarity in satellite time series between planted fields and other vegetated areas, which may include fallow fields in addition to non-agricultural areas. We used the same reference samples to compute class-wise accuracy metrics (overall accuracy, user accuracy/precision, producer accuracy/recall, true positive rate, and false positive rate) in Table 5.
The change estimates at the admin2 scale follow a similar trend as the admin1 (Tigray) scale. The majority of the area in each zone was estimated as stable non-crop followed by stable cropland. A relatively small amount of cropland gain and loss was estimated in each zone. However, the standard errors relative to the area estimates were much larger at the admin2 scale compared to those for Tigray due to smaller reference sample sizes. The standard error was highest for the change classes in Western Tigray. Figures 7(a)-11(a) provide the confusion matrices for each admin2 zone. Tables 7-11 report the accuracy metrics computed using the reference samples in each zone.
@@
Table 4: Change area estimates for Tigray and each zone (admin2 level) within Tigray (units are kha)
@@
Table 5: Accuracy metrics for four-class cropland change reference sample in Tigray. Overall accuracy is $0.77\pm 0.02$. See Tables 7-11 for admin2/zone-level metrics.

### v2

@@ 2.3.5 Post-processing
The University of Ghent and Mekelle University published a small ground-truth survey dataset of 161 fields surveyed in Tigray at the end of August 2021 [29]. The team of geographers visited six districts in Eastern and Southern Tigray (including Mekelle), spanning ecoregions with diverse biophysical and agro-ecological characteristics. Fields surveyed were typically near (hundreds of meters away from) main roads. Selection criteria for fields included rainfed farmland (no irrigation); no buildings, iron sheets, or trees on the land; and field size at least $30$ m $\times$ $30$ m, and no intercropping. Surveyors recorded point locations in the center of each field. Data for each field was recorded by observing and talking to farmers who were present on the land.
Surveyors recorded the crop type being grown in each field as well as the crop condition (e.g., good, medium, poor) and notes about the nearby farmlands. Of the 34 fields recorded as fallow, 20 were recorded as “fallow with weeds or grass” and 14 as “fallow (no vegetation at all)” (meaning the field had been plowed but not seeded). The 127 remaining cultivated fields included teff (33), maize (7), sorghum (5), other cereals such as wheat, barley, or millet (62), oil crops such as flax or niger seed (12), legumes such as beans or peas (6), and potato (2).
@@
To better understand the difficulty of distinguishing cultivated from fallow fields (which may resemble cultivated fields due to the presence of weeds or other vegetation) in satellite observations, we plotted the 12-month time series profiles for these ground-truth points in each of the 12 Sentinel-2 channels that were used for our model input. Figure 3 shows the NDVI time series and 14 shows the time series for all Sentinel-2 bands. Figure 14 shows that the time series for all bands is visually very similar for the cultivated and fallow fields surveyed. This observation, in addition to visual inspection of predicted crop maps in Tigray, suggested that our predicted crop maps may overestimate cultivated area in Tigray by classifying some fallow fields as cultivated.
@@
Figure 3: Left: 12-month time series of Sentinel-2 NDVI for ground-truth plots annotated as cultivated (“Crop”), fallow with weeds (“Fallow Weed”), or fallow with no vegetation at all (“Fallow None”). Solid lines depict the mean and shaded regions depict the standard deviation of the time series for all samples. Right: Visualizations of example ground-truth plots in PlanetScope September 2021 basemap. The plot labeled “Fallow with weeds or grass” has a similar appearance to fields labeled with crop types, while the field labeled “Fallow (no vegetation at all)” is easier to distinguish. Note that our map classification used Sentinel-2 inputs (left-most time series) while our response design used PlanetScope data for photointerpretation of reference sample labels (similar to those shown on the right).
@@
In Figure 3, some differentiation between cultivated and fallow fields can be seen in the NDVI time series: the peak of the NDVI time series appears substantially higher on average than the peak reached in fallow fields. This also makes intuitive sense: we would expect fallow fields to have lower biomass than cultivated fields, since the pattern of vegetation in fallow fields should be sparse and unmaintained. Thus, we hypothesized that for a given predicted cropland map, within the pixels predicted as cropland there will be a minority sub-population of pixels that reached lower peak NDVI than the majority of the pixels which may correspond to fallow fields (or other erroneous predictions). To identify and remove these potential false positives, we computed the mean and standard deviation of the peak NDVI value for all positive (cropland) predictions in a given map. We then identified which of those pixels had NDVI below $n$ standard deviations from the mean (i.e., $\mu-n\sigma$). We then re-classified those “anomalous” positive pixels to the non-crop class.
We evaluated the effect of using $n\in\{0,0.5,1.0,1.5,2.0,2.5,3.0,3.5,4.0\}$ for this post-processing algorithm on the true positive rate (TPR) versus false positive rate (FPR) for the ground-truth dataset and validation dataset for Tigray in 2021. Figure 4 shows a plot of these values. We found that a threshold of $n=3.5$ offered the optimal tradeoff between TPR and FPR for the ground-truth dataset. This post-processing threshold also increased the TPR on the validation dataset without significantly increasing the FPR. Thus for our final cropland maps in 2020 and 2021, we applied this post-processing algorithm with $n=3.5$ to reduce false positive predictions.
Figure 4: True positive rate (TPR) versus false positive rate (FPR) for range of thresholds used for post-processing algorithm, evaluated for the ground-truth dataset and validation dataset for Tigray in 2021.
2.4 Comparison to existing land cover maps
Some prior studies used existing global land cover maps to derive a cropland map for change assessments instead of creating a custom classification map as in our study. For example, Weldegebriel et al. [19] used the WorldCover 10 m/pixel global land cover map for 2020 [30] to assess changes in cultivated area between 2020-2021 in Tigray.
To assess the accuracy of our custom map classification compared to existing land cover maps, we evaluated multiple global land cover maps using our independent test set following the methodology described in Kerner et al. [31]. We only included maps that were created in 2017 or later to ensure temporal coherence: WorldCover [32, 33], Digital Earth Africa [34, 35], Copernicus [36, 37], Dynamic World [38], ASAP [39, 40], GLAD [41, 42], and Esri LULC [43, 44]. If multiple years were provided by a land cover product, we used the closest year to our reference data year.
@@ 2.5.1 Sampling and response design
We created a stratified random sample of $10\times 10$-meter pixels using the four strata in the 2020-2021 cropland change map classes. We drew 799 samples total (sampling rate 0.00015% in Tigray) using the following allocation to the four strata: 131 for “stable planted,” 404 for “stable not-planted,” 134 for “planted gain,” and 130 for “planted loss.” We pre-allocated 100 samples to each change stratum, then allocated the remaining samples approximately proportionally to the map area for each stratum. Pre-allocation helps reduce the uncertainties in area estimates for change classes resulting from low sample sizes [20].
Each point was labeled by the authors and other trained individuals. Annotators answered two questions for each point: “Does this point contain active cropland in 2020?” and “Does this point contain active cropland in 2021?” Each reference sample was independently labeled by at least two annotators. For samples that did not have a unanimous label agreement for both questions, we convened expert annotators as a group to discuss and decide the label to assign for each year. We describe our labeling procedure used to answer this question in detail in Section 2.5.2.
2.5.2 Satellite interpretation of reference points
The labels assigned to the reference samples described in Sections 2.3.4 and 2.5.1 have a significant impact on the area estimates and performance metrics derived from those samples. In this section, we detail our procedure for determining whether a given reference point for a given year should be labeled as containing active cropland or not. We provide several examples to illustrate how we determined these labels in A.
We used Collect Earth Online (CEO) as our primary labeling interface. We created a CEO project for each set of reference points and configured the project’s Imagery Options to include the 3-5 m/pixel Planet Monthly Mosaics, 10-20 m/pixel Sentinel-2 monthly composite (displaying the Agriculture band combination – a false-color composite of the SWIR-1 (B11), near-infrared (B8), and blue (B2) bands – by default), and the 0.5-1 m/pixel Mapbox Satellite layers. The Planet and Sentinel-2 layers allow users to view the imagery for a specific month/year, while the Mapbox layer is a basemap and the date cannot be configured. Thus, we instructed labelers to base their interpretation primarily on Planet and Sentinel-2 since the date of the Mapbox images was unknown – however, the high-resolution Mapbox basemap can still provide additional context for the coarser Planet and Sentinel-2 layers.
For points where additional context might be helpful, labelers were instructed to click “Download Plot KML” in the CEO interface and load the point in Google Earth Pro on their desktop computer. In Google Earth Pro, labelers used the “Historical Imagery” tool to view any high-resolution imagery (typically provided by Maxar or Airbus/CNES) available for the point. While high-resolution images during the growing season of the reference year were rarely available, high-resolution images acquired at any point (or multiple points) in time can provide additional helpful context for interpreting the coarser imagery in CEO (similar to the Mapbox basemap). In Google Earth Pro, labelers can also change the viewing angle of the point and use the terrain layer to better visualize the landscape in 3D – this is particularly helpful for points on steep slopes or ridges unlikely to host agriculture.
The majority of agricultural land in Tigray is used to grow teff, barley, wheat, maize, sorghum, finger millet, oats, and sesame [23]. These are row crops that are planted and harvested on an annual cycle, necessitating yearly tilling, sowing, and management. Thus, as in Skakun et al. [6], we defined cropland as “a piece of land that is sown/planted and harvestable at least once within the 12 months after the sowing/planting date.” This is also consistent with the Ethiopia Central Statistics Agency’s definition of annual/temporary crops as “crops, which are grown in less than a year’s time, sometimes only a few months with an objective to sow or replant again for additional production following the current harvest. Continuously grown crops planted in rotation are also considered as temporary crops since each is harvested and destroyed by ploughing in preparation for each successive crop” [23]. Our definition does not include permanent (perennial) crops, since these represent a minority of production in Tigray [23]. Our definition also does not include livestock or pasture.
Labelers typically stepped through the Planet Monthly Mosaic for each month spanning the crop calendar for Tigray (considering both the belg and meher seasons), zooming in/out and panning around to assess the point locally and compared to nearby fields or landscapes. They did the same for the Sentinel-2 layer and compared the Mapbox layer to the Planet and Sentinel-2 layers for context. If needed, they downloaded the point’s KML file and viewed historical imagery available in Google Earth Pro. Labelers looked for the following indicators of active cultivation of annual crops:
  • 1.
    Approximately rectilinear shapes of fields
  • 2.
    Relatively uniform color and texture within field boundary suggesting management
  • 3.
    Dark brown color during land preparation/planting months indicating plowing
  • 4.
    Changes in color and texture aligned with crop calendar: e.g., increase in greenness during months between sowing and vegetated phases, dark green color during peak vegetation month(s), or rapid changes from green to brown in harvest months
  • 5.
    Presence of harvest piles at the end of the growing season [46]
In contrast, fallow (inactive) fields tend to be characterized by:
  • 1.
    Lack of uniformity in color and texture
  • 2.
    Sparse, mottled, or un-managed appearance
  • 3.
    Brown, beige, or a lighter green color indicating the presence of bare soil, weeds, or residual plants
  • 4.
    Lack of significant change in color and texture throughout the growing season
While all of these characteristics may not be visible in the imagery for every point, labelers looked for a combination of these characteristics to decide the label for a point. We illustrate some of these characteristics in the examples provided in A.
2.5.3 Cultivated area in Tigray region and zones
@@
We used the reference samples to compute confusion matrices between the map class and reference labels, which were then used to compute map accuracy, area estimates and uncertainties, and investigate changes between 2020-2021 that could be seen visually in satellite images [20]. We computed area estimates for the four change classes representing transitions from 2020 to 2021 as well as annual cropland area estimates for 2020 and 2021 separately. For the annual area estimates, we used the reference labels from each year provided by the two-year change reference samples. We computed the annual and change area estimates for the Tigray region as well as within each of the five zones in Tigray. For each zone, we clipped the classification map to the zone boundary and filtered the reference samples to use only the samples within a given zone to estimate area. The code used to compute annual area estimates is available at https://github.com/nasaharvest/crop-mask/blob/165d01cc3114427fac59f28ec2eeee5c97de25fc/notebooks/crop_area_estimation.ipynb. The code used to compute change area estimates is available at https://github.com/nasaharvest/crop-mask/blob/359708379fb7f064e1248ff35daaf56e27eee97a/notebooks/ethiopia_tigray_change_area_estimation.ipynb. The reference samples can be accessed via Github following the paths recorded in the change area estimation notebook.
@@ 3.1 Map classification and accuracy assessment
Figure 5: Annual cropland and change maps overlaid with zone boundaries (solid lines) and 5 km buffer around conflict events (dashed lines).
Table 1 and Table 2 report the overall accuracy, precision (user’s accuracy), recall (producer’s accuracy), and F1 score for the cropland class in the 2020 and 2021 maps, respectively. We computed these metrics using the independent test sets labeled for each year described in Section 2.3.4. We found F1 scores of $0.78\pm 0.11$ for the 2021 map and $0.69\pm 0.11$ for 2020. Overall accuracy was similar between the two years: $0.83\pm 0.02$ in 2020 and $0.84\pm 0.02$ in 2021. There was a moderate difference in precision (user’s accuracy) between 2020 ($0.75\pm 0.04$) and 2021 ($0.79\pm 0.04$) and larger differences in recall (producer’s accuracy) between 2020 ($0.65\pm 0.03$) and 2021 ($0.76\pm 0.03$). This is likely due to differences in climatic factors between the two years that result in different types of errors made by the classifier. For example, differences in rainfall between the two years could affect vegetation patterns that make cultivated cropland easier or harder to distinguish from fallow land or other vegetation, or differences in cloud cover could affect the quality of the input time series in some areas. Unbalanced precision and recall scores between the two years may result in higher error rates when the maps are combined to form a change map, compared to maps with balanced scores.
@@
Table 1: Performance metrics for 2020 test set used for assessing accuracy of annual cropland map. We report metrics for other cropland and land cover maps on the same test set for comparison to our map (ordered by decreasing F1 score). Legend: OA = Overall Accuracy, PA = Producer’s Accuracy, UA = User’s Accuracy, TN = True Negatives, FP = False Positives, FN = False Negatives, TP = True Positives.
@@
Table 2: Performance metrics for 2021 test set used for assessing accuracy of annual cropland map. We report metrics for other cropland and land cover maps on the same test set for comparison to our map (ordered by decreasing F1 score). Legend: OA = Overall Accuracy, PA = Producer’s Accuracy, UA = User’s Accuracy, TN = True Negatives, FP = False Positives, FN = False Negatives, TP = True Positives.
Of the compared maps described in Section 2.4 (not including our map), we found that WorldCover had the highest performance metrics overall for both years. Compared to our map, the WorldCover map had a similar F1 score in 2020 but had significantly lower performance in 2021.
@@
Table 3: Annual area estimates for Tigray and each zone within Tigray (units are kha)
@@ 3.2 Annual area estimates
We report the annual cropland area estimates for 2020 and 2021 in Table 3. Figure 6 shows these estimates as a bar plot to illustrate the relative differences between each year. We found similar estimates for cultivated area in Tigray in 2020 and 2021, with a possible small net increase: $1,132\pm 133$ kha in 2020 and $1,217\pm 132$ kha in 2021 (1 kha = 1,000 ha). We found a similar trend for each zone, with the exception of the Central zone which showed a possible small net decrease. This net decrease in Central Tigray may be due to a higher concentration of conflict events recorded in the Central zone during the growing season [19]. Figure 7 shows the confusion matrices between the map and reference samples for Tigray. We provide the confusion matrix for each zone in Figures 15-19. All confusion matrices are expressed in terms of proportion of mapped area based on good practice recommendations [20].
Figure 6: Annual area estimates for Tigray region and each zone within Tigray (units are kha)
@@ 3.3 Change area estimates
Table 4 summarizes the area estimates for the four transition classes. We found that cropland gain accounted for $70-160$ kha while cropland loss accounted for $19-79$ kha in Tigray. The stable transition classes accounted for the majority of the area in Tigray: $1,033-1,293$ kha of stable cropland and $3,798-4,054$ kha of stable non-cropland. Figure 7(a) shows the confusion matrix between the reference samples (stratified by the four change class strata) and the classified change map, expressed in terms of mapped area proportions. For Tigray, there was good agreement between the map and reference samples for the stable not planted class: of the area mapped as stable not planted, only a small fraction of the mapped area should have been mapped as stable planted according to the reference sample. There was also good agreement between the map and reference samples for the stable planted class, though approximately one third of the area mapped as stable planted should have been mapped as stable not planted according to the reference sample. This is likely due to the similarity in satellite time series between planted fields and other vegetated areas, which may include fallow fields in addition to non-agricultural areas. We used the same reference samples to compute class-wise accuracy metrics (overall accuracy, user accuracy/precision, producer accuracy/recall, true positive rate, and false positive rate) in Table 5.
The change estimates at the zone scale follow a similar trend as the region (Tigray) scale. The majority of the area in each zone was estimated as stable non-crop followed by stable cropland. A relatively small amount of cropland gain and loss was estimated in each zone. However, the standard errors relative to the area estimates were much larger at the zone scale compared to those for Tigray region due to smaller reference sample sizes. The standard error was highest for the change classes in Western Tigray. Figures 15(a)-19(a) provide the confusion matrices for each zone. Tables 7-11 report the accuracy metrics computed using the reference samples in each zone.
While we provided the full-resolution cropland and cropland change maps as part of this study for transparency and reproducibility of our area estimates, we emphasize that their intended use is for stratification in the area estimation procedure (see Figure 2(b)). The change map has limited accuracy as shown in Figures 15-19 and is not intended for visual interpretation of changes or other uses outside of the sample-based area estimation described in this paper. We also emphasize that inaccuracies in the map do not signify inaccuracies in the area estimates, since such errors are accounted for by the use of an unbiased estimator [45].
@@
Table 4: Change area estimates for Tigray region and each zone within Tigray (units are kha)
@@
Table 5: Accuracy metrics for four-class cropland change reference sample in Tigray. Overall accuracy is $0.77\pm 0.02$. See Tables 7-11 for zone-level metrics. Legend: OA = Overall Accuracy, PA = Producer’s Accuracy, UA = User’s Accuracy, TN = True Negatives, FP = False Positives, FN = False Negatives, TP = True Positives.